With more rain expected in the coming days/weeks comes the danger of flood, erosion and silting damage, as well as rock and mud slides, trees blowing down, stumps breaking loose and rolling down hillsides, etc. Sludge was already on the road this morning, as seen in the picture below.
|Sludge from the severe burns above
the road around mile 133 on Swift Trail (LBTO/P. Hartley)
|Rain on the SW side of Mt Graham
(LBTO – roofcam)
|The inflatable bladder installed on the roof (red tube) between the shutter doors and the central plenum (LBTO/J. Riedl)|
|Animation of the Frye Fire evolution from June 13 to July 9
(NIROPS – Google Earth – LBTO)
The Frye Fire
|Back burns north of the observatory (LBTO/S. Allanson)|
Telescope operator and instrument specialist leave the observatory. The fire comes closer…
A fire storm gets close to LBTO and within feet of the VATT. A walk around all of the buildings shows that they have taken some heat, but none appear to have been damaged. Power remains on…
|Through the kitchen window
|A DC10 dumping retardant
We are able to relieve the mountain manager (Kevin down, Pat up).
|A pink SMT and a still standing VATT on the morning after the fire storm (LBTO/P. Hartley)|
|Still burning not too far from the site (LBTO/roofcam)|
Now, we have a skeleton crew of 2 people…
Down to a crew of 1.
|10mn of rain give a short-lived hope that real rain will come… (LBTO/J. Riedl)|
MGIO water truck deliveries have been shut down after one delivery. Site running on generator.
|Still burning west of MGIO, though the observatories are safe (LBTO/roofcam)|
No traffic authorized…
|Wind revives close-by smoldering areas and ashes are flying. No harm done as winds calm down as the night progresses. (LBTO/roofcam)|
|On swift trail… (LBTO/J. Riedl)|
|Winds are in the right direction, blowing the smoke away.
We can open the vent door to cool down the chamber! (LBTO)
A short visit by VATT staff of their telescope shows no damage on the telescope. A big relief!
|VATT is unscathed! (VATT)|
Crew shift change and a crew of 4 at LBTO. Still waiting for a fuel delivery as the site continues to run on generator.
|Morning smoke on a nice clear sky… (LBTO)|
|On the way up (unpaved section of
Swift Trail) (LBTO/M Wagner)
|Meanwhile, view from the LBTO roof… (LBTO/roofcam)|
No traffic (!)
|The following morning (LBTO/M. Wagner)|
Early departure from base camp (5:30am) with W. Seifert and LBTO staff. LUCI1 servicing mission will be able to start with only a one-day delay.
|On the way up (LBTO/J. Riedl)|
Seven people are at LBTO!
|Clouds and smoke… Still hoping for rain to be able to open before the end of the semester (5 nights left)
|Still burning SW of the observatory… (LBTO)|
|Smoke is still present and clouds are rolling in. We will be closed tonight. With more clouds and thunderstorms to come, the chance to open before the beginning of the shutdown is very slim! (LBTO)|
First day of shutdown… Crew swap with the iLocater team leaving (happy with what they achieved) and the LUCI Aux Cryostat team coming up.